# From smooth dynamical twists to twistors of quantum groupoids

### v1

@@ 3. From dynamical twists to twistors of quantum groupoids
We start by introducing a quantum groupoid arising from a Lie algebra pair $({\mathfrak{g}},{\mathfrak{l}})$. Let $L\subset G$ be two Lie groups integrating ${\mathfrak{l}}$ and ${\mathfrak{g}}$, respectively. For an $L$-invariant open submanifold $V\subset{\mathfrak{l}}^{\ast}$, the product $V\times G$ admits a natural left $G$-action defined by
@@
In [Gutt], Gutt constructed an explicit star product on the cotangent bundle $T^{\ast}L\cong{\mathfrak{l}}^{\ast}\times L$, which realizes deformation quantization of the Poisson structure on
@@
coming from the canonical symplectic structure on the cotangent bundle. This star product, which we call Gutt star product, is given by
@@ 3.1. Main result
Composing with the coproduct $\Delta$ on ${\mathcal{D}}\llbracket\hbar\rrbracket$, we obtain
@@
$\in\left({\mathcal{D}}\llbracket\hbar\rrbracket\otimes_{R}{\mathcal{D}}\llbracket\hbar\rrbracket\right)\otimes\left(U({\mathfrak{g}})\otimes U({\mathfrak{g}})\right)\cong{\mathcal{H}}\otimes_{R}{\mathcal{H}},$ (3.2)
@@
    When the Lie subalgebra ${\mathfrak{l}}$ is abelian, the Gutt twistor $\Theta_{\mathrm{Gutt}}=\Theta_{0}=\sum_{i}\hbar\frac{\partial}{\partial\lambda_{i}}\otimes l_{i}$ and the PBW star product $\ast_{{\mathrm{PBW}}}$ coincide with the usual multiplication of functions. Applying Theorem 3.1 and Corollary 3.2 to this case, we recover the construction of twistors from dynamical twists and dynamical quantum groupoids in [Xu-R-matrices].
@@
    Suppose that $F=\sum_{r,s}\hbar^{p}b_{p,rs}\otimes U_{p,r}\otimes U_{p,s}\in C^{\infty}(V)\llbracket\hbar\rrbracket\otimes U({\mathfrak{g}})\otimes U({\mathfrak{g}})$ is a dynamical twist. Then its classical limit $\theta=\sum_{r,s}b_{1,rs}\otimes(U_{1,r}\otimes U_{1,s}-\otimes U_{1,s}\otimes U_{1,r})$ is a triangular classical dynamical $r$-matrice which satisfies the classical dynamical Yang-Baxter equation (CDYBE)[Xu-Nonabelian]. On the other hand, consider the associated twistor
@@ 3.2. More facts about the Gutt twistor
$[\Delta(\lambda_{i}\ast_{{\mathrm{PBW}}}\,+\,\hbar l_{i})\,,\overline{F\ast_{{\mathrm{PBW}}}}]=0\in{\mathcal{H}}\otimes_{R}{\mathcal{H}}.$ (3.4)
@@
We first note that two differential operators $D_{1}=D_{2}\in{\mathcal{D}}$ if and only if $D_{1}g=D_{2}g$ for all polynomial functions $g\in S({\mathfrak{l}})$. Hence, for all $g_{1}\in S({\mathfrak{l}}),g_{2},g_{3}\in C^{\infty}(V\times G)$,
@@
Moreover, since $(\epsilon\otimes_{R}{\mathrm{id}})\Theta_{\mathrm{Gutt}}=({\mathrm{id}}\otimes_{R}\epsilon)\Theta_{\mathrm{Gutt}}=1\otimes_{R}1$, it follows that ${\mathcal{F}}$ satisfies Equation (2.6) if and only if $F$ satisfies Equation (2.11). Therefore, $F$ is a dynamical twist if and only if ${\mathcal{F}}=\overline{F\ast_{{\mathrm{PBW}}}}\cdot\Theta_{\mathrm{Gutt}}$ is a twistor of the quantum groupoid ${\mathcal{H}}$.
@@ 4. The Enriquez-Etingof formal dynamical twist
Various constructions of equivariant star products on Poisson homogeneous spaces from dynamical twists have been well-studied (see [Donin-Mudrov-1, Enriquez-Etingof2, Enriquez-Etingof-Marshall, Karolinsky-Muzykin-Stolin-Tarasov, Karolinsky-Stolin-Tarasov-2]). As an application of Theorem 3.1, we provide an approach to constructing such equivariant star products from the quantum groupoid point of view.
@@
where ${\mathfrak{u}}_{+}$ and ${\mathfrak{u}}_{-}$ are Lie subalgebras of ${\mathfrak{g}}$ satisfying $[{\mathfrak{l}}\,,{\mathfrak{u}}_{+}]\subset{\mathfrak{u}}_{+}$ and $[{\mathfrak{l}}\,,{\mathfrak{u}}_{-}]\subset{\mathfrak{u}}_{-}$.
@@
for all $u,v\in{\mathfrak{u}}$. By fixing an isomorphism ${\mathfrak{u}}^{\ast}\simeq{\mathfrak{u}}$ of ${\mathbb{K}}$-vector spaces, we may view $\omega(\lambda)$ as an element in $\operatorname{End}({\mathfrak{u}})$.
A polarized Lie algebra ${\mathfrak{g}}$ is called nondegenerate if there exists some $\lambda\in{\mathfrak{l}}^{\ast}$ such that $\omega(\lambda)$ is nondegenerate, which is equivalent to that $D(\lambda):=\det\omega(\lambda)\neq 0$. Let $\widehat{S}({\mathfrak{l}})[1/D]$ be the localization of the formal completion of $S({\mathfrak{l}})$ with respect to $D$. A deformation quantization of $\widehat{S}({\mathfrak{l}})[1/D]$ is provided by a completed associative ${\mathbb{K}}\llbracket\hbar\rrbracket$-algebra $\widehat{S}({\mathfrak{l}})[1/D]_{\hbar}$ [Enriquez-Etingof2]. Moreover, it is proved that there exists a unique ${\mathfrak{l}}$-invariant element
@@
Assume that the closed submanifold $W\subset V$ consisting of characters of 1-dimensional trivial representations of ${\mathfrak{l}}$ is nonempty. Let $G\supset L$ be two Lie groups integrating ${\mathfrak{g}}$ and ${\mathfrak{l}}$, respectively. Consider the space of $W$-parameterized family of smooth differential operators on the homogeneous space $G/L$

### v2

@@ 3. From dynamical twists to twistors of quantum groupoids
We start by introducing a quantum groupoid arising from a Lie algebra pair $({\mathfrak{g}},{\mathfrak{l}})$. Let $G$ be a Lie group integrating $\mathfrak{g}$} and $L\subset G$ be a Lie subgroup integrating ${\mathfrak{l}}\subset{\mathfrak{g}}$. For an $L$-invariant open submanifold $V\subset{\mathfrak{l}}^{\ast}$, the product $V\times G$ admits a natural left $G$-action defined by
@@
In [Gutt], Gutt constructed an explicit star product on the cotangent bundle $T^{\ast}L\cong{\mathfrak{l}}^{\ast}\times L$, which realizes deformation quantization of the Poisson structure
@@
coming from the canonical symplectic structure on the cotangent bundle. A normal ordering version of this star product, which we also call Gutt star product, is given in [Xu-Nonabelian]
@@ 3.1. Main result
[…]
@@
$\in\left({\mathcal{D}}\llbracket\hbar\rrbracket\otimes_{R}{\mathcal{D}}\llbracket\hbar\rrbracket\right)\otimes\left(U({\mathfrak{g}})\otimes U({\mathfrak{g}})\right)\cong{\mathcal{H}}\otimes_{R}{\mathcal{H}}.$ (3.2)
@@
    When the Lie subalgebra $\mathfrak{l}$} is abelian, the Gutt twistor $\Theta_{\mathrm{Gutt}}=\Theta_{0}=\exp(\sum_{i}\hbar\frac{\partial}{\partial\lambda_{i}}\otimes l_{i})$ and the PBW star product $\ast_{{\mathrm{PBW}}}$ coincide with the usual multiplication of functions. Applying Theorem 3.1 and Corollary 3.2 to this case, we recover the construction of twistors and dynamical quantum groupoids from smooth dynamical twists in [Xu-R-matrices].
@@
    Suppose that $F=\sum_{p=0}^{\infty}\sum_{r,s}\hbar^{p}b_{p,rs}\otimes U_{p,r}\otimes U_{p,s}\in C^{\infty}(V)\llbracket\hbar\rrbracket\otimes U({\mathfrak{g}})\otimes U({\mathfrak{g}})$ is a smooth dynamical twist. Then its classical limit $\theta=\sum_{r,s}b_{1,rs}\otimes(U_{1,r}\otimes U_{1,s}-\otimes U_{1,s}\otimes U_{1,r})$ is a triangular classical dynamical $r$-matrice which satisfies the classical dynamical Yang-Baxter equation (CDYBE)[Xu-Nonabelian]. On the other hand, consider the associated twistor
@@ 3.2. More facts about the Gutt twistor
$\Delta(\lambda_{i}\ast_{{\mathrm{PBW}}}\,+\,\hbar l_{i})\,,\overline{F\ast_{{\mathrm{PBW}}}}]=0\in{\mathcal{H}}\otimes_{R}{\mathcal{H}}.$ (3.4)
@@
We first note that two differential operators $D_{1}=D_{2}\in{\mathcal{D}}$ if and only if $D_{1}g=D_{2}g$ for all polynomial functions $g\in S({\mathfrak{l}})$. Hence, the calculation below, which holds for all $g_{1}\in S({\mathfrak{l}}),g_{2},g_{3}\in C^{\infty}(V\times G)$, verifies the statement:
@@
Moreover, since $(\epsilon\otimes_{R}{\mathrm{id}})\Theta_{\mathrm{Gutt}}=({\mathrm{id}}\otimes_{R}\epsilon)\Theta_{\mathrm{Gutt}}=1\otimes_{R}1$, it follows that $\mathcal{F}$} satisfies Equation (2.6) if and only if $F$ satisfies Equation (2.11). Therefore, $F$ is a smooth dynamical twist if and only if ${\mathcal{F}}=\overline{F\ast_{{\mathrm{PBW}}}}\cdot\Theta_{\mathrm{Gutt}}$ is a twistor of the quantum groupoid $\mathcal{H}$}.
@@ 4. The Enriquez-Etingof formal dynamical twist
Various constructions of equivariant star products on Poisson homogeneous spaces from dynamical twists have been well-studied (see [Donin-Mudrov-1, Enriquez-Etingof2, Enriquez-Etingof-Marshall, Karolinsky-Muzykin-Stolin-Tarasov, Karolinsky-Stolin-Tarasov-2]). As an application of Theorem 3.1, we provide an approach to construct such equivariant star products from the quantum groupoid point of view.
@@
where $\mathfrak{u}$$_{+}$ and $\mathfrak{u}$$_{-}$ are Lie subalgebras of $\mathfrak{g}$} satisfying ${\mathfrak{l}}\,,{\mathfrak{u}}_{+}]\subset{\mathfrak{u}}_{+}$ and ${\mathfrak{l}}\,,{\mathfrak{u}}_{-}]\subset{\mathfrak{u}}_{-}$.
@@
for all $u,v\in{\mathfrak{u}}$. By fixing an isomorphism ${\mathfrak{u}}^{\ast}\cong{\mathfrak{u}}$ of $\mathbb{K}$}-vector spaces, we may view $\omega(\lambda)$ as an element in $\operatorname{End}({\mathfrak{u}})$.
[…]
@@
Assume that the closed submanifold $W\subset V$ consisting of characters of 1-dimensional trivial representations of $\mathfrak{l}$} is nonempty. Let $G$ be a Lie group integrating $\mathfrak{g}$} and $L\subset G$ be a Lie subgroup integrating ${\mathfrak{l}}\subset{\mathfrak{g}}$. Consider the space of $W$-parameterized family of smooth differential operators on the homogeneous space $G/L$